### v1

Abstract
The electrification and automation of mobility are reshaping how cities operate on-demand transport systems. Managing Electric Autonomous Mobility-on-Demand (EAMoD) fleets effectively requires coordinating dispatch, rebalancing, and charging decisions under multiple uncertainties, including travel demand, travel time, energy consumption, and charger availability. We address this challenge with a combined stochastic and robust model predictive control (MPC) framework. The framework integrates spatio-temporal Bayesian neural network forecasts with a multi-stage stochastic optimization model, formulated as a large-scale mixed-integer linear program. To ensure real-time applicability, we develop a tailored Nested Benders Decomposition that exploits the scenario tree structure and enables efficient parallelized solution. Stochastic optimization is employed to anticipate demand and infrastructure variability, while robust constraints on energy consumption and travel times safeguard feasibility under worst-case realizations. We evaluate the framework using high-fidelity simulations of San Francisco and Chicago. Compared with deterministic, reactive, and robust baselines, the combined stochastic and robust approach reduces median passenger waiting times by up to 36% and 95th-percentile delays by nearly 20%, while also lowering rebalancing distance by 27% and electricity costs by more than 35%. We also conduct a sensitivity analysis of battery size and vehicle efficiency, finding that energy-efficient vehicles maintain stable performance even with small batteries, whereas less efficient vehicles require larger batteries and greater infrastructure support. Our results emphasize the importance of jointly optimizing predictive control, vehicle capabilities, and infrastructure planning to enable scalable, cost-efficient EAMoD operations.
keywords:
[…]
1 Introduction
As cities worldwide face rapid urban growth, environmental challenges, and evolving societal demands, transportation systems play a critical role in achieving sustainable urban development. Expanding urban populations exacerbate issues such as traffic congestion, pollution, aging infrastructure, and socioeconomic inequality, threatening the sustainability and livability of urban areas [1, 2]. In response, the concept of smart cities has emerged, leveraging advanced technologies such as Intelligent Transportation Systems (ITS). ITS integrate real-time data, connected infrastructure, and autonomous vehicles to enhance urban management, optimize mobility, and promote sustainable transportation solutions [3, 4, 5]. A promising development within this domain is Electric Autonomous Mobility-on-Demand (EAMoD) systems. These systems consist of centrally coordinated fleets of electric autonomous vehicles that provide on-demand transport services [6]. EAMoD systems have the potential to reduce urban emissions, improve operational efficiency, and accelerate the adoption of electric vehicles (EVs). Unlike fixed-route public transport, EAMoD offers dynamic routing and centralized coordination, enabling adaptation to real-time demand [7].
Optimizing EAMoD systems introduces several challenges, particularly in addressing demand-supply imbalances resulting from spatial and temporal variations in travel demand. Numerous studies have proposed various vehicle allocation control methods to mitigate these imbalances. For instance, [8] introduced a Model Predictive Control (MPC) approach that demonstrated promising results; however, its scalability was limited due to the need to model the entire road network. To enhance scalability, many studies have adopted discretizations strategies, dividing the city into larger areas, often referred to as stations. An example of this approach is presented in [9], where a queue-theoretical model was employed. An alternative and effective model-free approach is the "+1 method," which directs vehicles to the most recent service request location, as demonstrated by [10]. For MPC methods, incorporating uncertainties into the control algorithm has proven important. A stochastic MPC framework combined with a Long Short-Term Memory (LSTM) neural network for travel demand prediction, leveraging Sample Average Approximation (SAA) to model uncertainty, was proposed by [11]. Similarly, [12] introduced a chance-constrained approach using Gaussian Process Regression (GPR) for demand prediction, while [13] proposed a data-driven distributionally robust optimization method. Despite these advancements, many existing strategies often overlook the complexities associated with electric vehicles (EVs), including limited battery capacities and charging requirements, as highlighted in the review by [7].
To address this gap, several studies have explored charge scheduling strategies in EAMoD. Charging strategies such as plug-in versus battery swapping have been investigated to optimize fleet availability and operational efficiency in EAMoD systems [14]. The placement of charging stations using stochastic programming and considering different uncertainties was studied in [15]. The charging dynamics based on data from an EV taxi fleet were examined in [16]. Fleet sizing and charging station placement considering endogenous congestion were analyzed in [17]. While these studies address important aspects of EAMoD fleets, the challenging task of simultaneously considering charge scheduling and vehicle allocation is not explored. Some of these aspects are considered in [18], where a distributionally robust approach was proposed to account for uncertainties in demand and supply. An ARIMA model was used for travel demand prediction, and the uncertainty set was constructed using bootstrapping. Different reinforcement learning algorithms have also been proposed. A queue-theoretical deep reinforcement learning method incorporating routing, battery charging, and dynamic pricing was introduced by [19]. Various reinforcement learning strategies are compared in the survey by [20]. However, a comprehensive approach that simultaneously addresses both charging scheduling and vehicle allocation while accounting for uncertainties in travel demand, charging availability, energy consumption and travel time remains unexplored.
A common approach to handling uncertainties in control and optimization is to construct scenario trees. A scenario tree provides a structured representation of the evolution of uncertainty over time in a multi-stage decision process. Each node in the tree corresponds to a possible future state, while the branches represent different ways in which uncertainty can unfold, along with their associated probabilities. It has been demonstrated that increasing the number of scenarios in the tree improves the accuracy of the uncertainty representation. However, it should be noted that an increase in the number of scenarios also results in an increase in the computational complexity of the optimization problem [21]. In order to address this issue, a range of scenario reduction techniques have been developed with the objective of constructing a reduced scenario tree that closely approximates the original one. These methods typically rely on a metric that quantifies the difference between the original and reduced trees, which can be used both to guide the reduction procedure and to evaluate its quality. A prevalent methodology in this field is moment matching, whereby the reduced scenario tree is constructed in order to preserve the first few moments (e.g., mean, variance, skewness, and kurtosis) of the original stochastic process [22]. This technique has been employed in several studies to generate scenario trees by matching the first four moments [23, 24]. However, it has been shown that two probability distributions with matching moments can still be significantly different [25]. An alternative metric is the Wasserstein distance, which quantifies the similarity between two probability distributions and has been used for scenario reduction [26]. However, the Wasserstein distance does not consider the filtration structure of the scenario tree, meaning it does not account for the temporal correlations between different stages. To address this, the Nested distance, also referred to as process distance, was introduced in [27]. The Nested distance explicitly accounts for the tree structure, preserving the temporal dependencies of the stochastic process. Although computing the Nested distance between two scenario trees is relatively straightforward, finding an optimally reduced scenario tree that minimizes this distance is computationally challenging. One approach to scenario reduction based on the Nested distance involves multistage k-means clustering, as proposed in [28]. However, this method still requires solving multiple large-scale linear programs (LPs), making it computationally demanding. A comprehensive evaluation of different scenario reduction techniques using the Nested distance is provided in [29], which highlights that methods minimizing the Nested distance tend to be computationally expensive. Another challenge is reducing scenario trees that incorporate multiple sources of uncertainty. In this paper, we propose a clustering-based technique for scenario tree reduction that effectively handles multiple sources of uncertainty.
Despite advances in scenario reduction, solving multi-stage stochastic programs (SPs) remains computationally challenging, particularly as the number of scenarios and the planning horizon increase. To address this, a wide range of decomposition techniques have been developed. The classical Benders decomposition method [30] laid the foundation for solving mixed integer linear programs (MILPs) with complicating variables, and was extended into the L-shaped method  [31] for two-stage SPs. Subsequent work introduced regularized variants [32] and stochastic decomposition methods [33, 34] inspired by stochastic quasi-gradient approaches [35] and stochastic approximation [36]. For multistage problems, nested Benders decomposition (NBD) [37] and multistage stochastic decomposition [38] were developed to extend these ideas. A major advancement came with sampling-based methods such as stochastic dual dynamic programming (SDDP) [39], and its subsequent enhancements [40, 41], as well as the stochastic dual dynamic integer programming (SDDiP) approach [42]. Alternative strategies include abridged nested Benders [43] and cutting-plane or partial sampling methods [44]. Recent approaches like stochastic dynamic linear programming [45] further illustrate the diversity of solution strategies. SDDP and SDDiP are especially effective for convex or mixed-integer multistage problems but rely on the assumption of stagewise independence of uncertainties, which limits their ability to capture temporal correlations. This assumption is often violated in real-world applications such as EAMoD systems, where travel demand, energy consumption, and charger availability exhibit strong temporal dependencies. To address these challenges, we propose a tailored nested Benders decomposition method that leverages the problem structure and captures temporally correlated, multivariate uncertainties.
In transportation research, variants of Benders decomposition has been successfully applied to a variety of different problems. A stabilized Benders algorithm has been used to solve a two-stage stochastic program for locating recharging stations in one-way electric car-sharing systems under demand uncertainty, demonstrating scalability to hundreds of large-scale demand scenarios [46]. An exact Benders decomposition embedded in a branch-and-cut framework has been developed to strategically locate mini-depots for last-mile crowd-shipping, achieving significant computational speedups and cost savings [47]. A hybrid approach combining Benders decomposition with NSGA-II has been proposed for the joint optimization of EV charging station placement and pricing, enabling efficient infrastructure deployment while balancing conflicting objectives such as consumer pricing and operator profitability [48].
The fleet composition, including the fleet size and individual vehicle attributes such as battery size and energy efficiency, affects how the fleet is optimized, the service quality provided, and the operational cost. For example, a less energy-efficient vehicle would need to charge more often compared to an efficient vehicle. Alternatively, the less energy-efficient vehicle could have a larger battery to compensate for its inefficiency, but this would increase costs and have a higher environmental impact. Few studies have incorporated fleet composition into the evaluation of EAMoD fleet performance. The effects of battery size and fleet size have been explored in [49]. However, fast charging and varying energy efficiency were not considered. In this paper, we address this research gap by investigating both battery size (40, 70, and 100 kWh), energy efficiency (efficient, medium efficient, inefficient), fleet size and charger availability with realistic DC fast charging dynamics [50].
These challenges, namely the integration of charge scheduling and vehicle allocation under uncertainty, the inclusion of heterogeneous fleet characteristics, and the construction of scalable scenario trees that capture multiple sources of uncertainty, remain largely unaddressed in the current literature. In this paper, we propose a novel framework for optimizing EAMoD systems by integrating predictive modeling with robust and stochastic optimization techniques to handle uncertainties in travel demand, charger availability, energy consumption, and travel time. Specifically, we adopt a robust approach to manage uncertainties in energy consumption and travel time, ensuring vehicles remain adequately charged and punctual to accommodate passenger pick-ups. For travel demand and charger availability, we optimize based on scenario MPC. To predict travel demand, we utilize a Bayesian Neural Network (BNN), which captures spatial and temporal dependencies between different areas of the city while providing a probabilistic framework to quantify uncertainty. For the BNN we follow the same structure as proposed in [51]. Based on the BNN predictions, scenarios are generated, and a reduced scenario tree is constructed using clustering techniques. The resulting optimization problem is formulated as a large-scale Mixed-Integer Linear Program (MILP), which is computationally challenging to solve. To address this, we propose a Nested Benders Decomposition (NBD) algorithm, enabling efficient real-time solutions. An overview of how the algorithm makes decisions based on data is presented in Fig. 1. The BNN model is benchmarked using taxi demand data from San Francisco and Chicago to demonstrate its predictive capabilities. Finally, the proposed optimization algorithm is rigorously tested within the high-fidelity transportation simulation framework AMoDeus [52].
The remainder of this work is organized as follows. Section 2 presents the model, the objective function, and the complete formulation of the multi-stage robust and stochastic MPC. The scenario generation and scenario reduction algorithm are introduced in Section 3. For scenario generation, a BNN is used to predict the uncertainty variables along with confidence estimation. Section 4 describes how the NBD is applied to the optimization problem. Finally, Section 5 presents the case studies used to benchmark the performance of the proposed algorithm. EAMoD transport simulations for San Francisco and Chicago are conducted under three different electricity price schemes, as well as for varying battery sizes and vehicle energy efficiencies. The section concludes with a summary of findings and directions for future research.
Historical data (travel demand, charger availability, energy consumption, travel time) + current system stateBNN: Bayesian Neural Network (predict demand, charging)Scenario Tree Generation (multi-stage forecasting)Scenario Tree Reduction (ScenTree algorithm)Prediction and Scenario Tree DetailsScenario Model Predictive Control (SMPC)Nested Benders Decomposition (NBD)Fleet and charging decisions
Figure 1: Overview of the scenario-based model predictive control (SMPC) framework using a Bayesian Neural Network (BNN), scenario tree generation and reduction, and Nested Benders Decomposition (NBD). Historical data is first used to train the BNN, which provides predictions of future demand and charging availability. These predictions are used to generate a multi-stage scenario tree, which is then reduced using the ScenTree algorithm to obtain a computationally tractable representation. The reduced scenario tree is used within the SMPC, which is solved using NBD to derive optimal fleet rebalancing and charging decisions.
2 EAMoD vehicle movement modeling
The model presented in this section is designed for centralized control of electric mobility-on-demand systems, focusing on optimizing vehicle movement and bi-directional charging scheduling to minimize costs while maximizing service completion levels. These high-level decisions serve as references for low-level controllers, which make real-time adjustments. Central to our model is its predictive nature across time, space, and energy dimensions. By exploiting its predictive capabilities, we aim to make optimal decisions in the current time step based on expected future states evolving in time and space. The model is tailored for receding horizon optimization, ensuring adaptability to changing conditions.
Our model is comprised of two integral components. First, we build upon the network model previously established in [12] by incorporating electric vehicles (EVs) into the framework. This enhancement requires adding several key features specific to EVs. First, charging stations to accommodate the recharging of vehicles, second the limited range of EVs necessitating strategic planning around vehicle deployment, and third vehicle-to-grid (V2G) capabilities, which allow for the interaction between the vehicle’s battery and the power grid for energy exchange. Secondly, we introduce uncertainties into the model, encompassing aspects such as demand variability, charging station availability, travel time, and energy consumption rates.
2.1 Model Notation
The Electric Autonomous Mobility on Demand (EAMoD) framework represents the city as comprising $n$ distinct stations or areas, enumerated as $\mathcal{N}=[1,\ldots,n]$. Stations are identified by applying a K-means clustering algorithm to travel data. The chargers, positioned at each station’s centroid, are publicly accessible and the charger occupancy is assumed to follow a non-stationary Poisson process. This simplification avoids the complexities of optimizing charger locations, for which existing literature provides insights [53, 54]. Travel is characterized by origin-destination pairs, with the origin denoted by index $i\in\mathcal{N}$ and the destination by index $j\in\mathcal{N}$. The state of charge (SoC) of vehicles is categorized into a set $\mathcal{B}=[0,0.01,\ldots,0.99,1]$, defining each SoC level by the index $b$. Time is discretized into $\mathcal{T}=[1,2,\ldots,T]$, where $t$ is the time index and $T$ the prediction horizon. A summary of different variables used in the model is provided in Table 1.
Table 1: Summary of all indices, state variables, decision variables, slack variables, and parameters used in the model.
Category Symbol Description
Indices
$i\in\mathcal{N}$ Origin of travel
$j\in\mathcal{N}$ Destination of travel
$b\in\mathcal{B}$ SoC
$t\in\mathcal{T}$ Time
$\xi\in\mathcal{S}$ Scenario sample
State Variables
$\lambda_{ij}\in\mathbb{N}$ Number of passengers traveling from $i$ to $j$
$\phi_{ib}\in\mathbb{N}$ Initial number of vehicles at station $i$ with SoC $b$
$\text{k}^{\text{c}}_{it}\in\mathbb{N}$ Number of available chargers at station $i$ at time $t$
Decision Variables
$x_{ibt}^{\text{c}}\in\mathbb{N}$ Vehicles charged at station $i$ with SoC $b$
$x_{ibt}^{\text{d}}\in\mathbb{N}$ Vehicles used for V2G at station $i$ with SoC $b$
$x_{ijbt}^{\text{p}}\in\mathbb{N}$ Vehicles transporting passengers from $i$ to $j$ with SoC $b$ at time $t$
$x_{ijbt}^{\text{r}}\in\mathbb{N}$ Vehicles rebalancing from $i$ to $j$ with SoC $b$ at time $t$
Slack Variables
$s_{ijt}\in\mathbb{N}$ Customers not served at station $i$ wishing to travel to $j$ at time $t$
Parameters
$\Delta b^{\text{c}}$ SoC charge in one time interval
$\Delta\text{d}$ SoC discharge to the grid in one time interval
$\text{e}^{\text{b}}$ Battery capacity [kWh]
$e_{ijt}$ Energy consumption for travel from $i$ to $j$ at time $t$
$P^{\text{c}}_{b}$ Charging power in kWh for different SoC levels $b$
$P^{\text{d}}$ V2G transfer power in kWh
$\Delta t_{ijt}$ Travel time between station $i$ and $j$ at time $t$
$\Delta b_{ijt}$ Energy consumption between station $i$ and $j$ at time $t$
$b_{ij}^{\text{min}}$ Minimum required SoC to drive a customer from station $i$ and $j$ at time $t$
$s$ Index for different scenarios
$S_{t}$ Set of scenarios at stage $t$
$b_{\text{max}}$ Max SoC threshold that is penalized in the terminal cost
$q_{\text{max}}$ Maximum penalty in terminal cost
2.2 Energy, Space and Time Network Model
To accurately represent the dynamic environment in which vehicles operate, we introduce an Energy, Space, and Time Network Model. This model forecasts the number of vehicles at the different stations in different time intervals, and their SoC levels, providing a comprehensive understanding of fleet operations (see Fig. 2). Rather than tracking individual vehicle trajectories, the model aggregates vehicles within each station based on identical states. This aggregation enhances computational tractability, enabling efficient analysis of fleet distribution and status across the network while maintaining manageable complexity. The stations are assumed to be fully interconnected, allowing unrestricted travel between any two locations. Consequently, the road network is modeled as a complete graph, ensuring total connectivity among all stations. In this framework, the travel time and energy consumption between any two stations $i$ and $j$ at time $t$ are denoted as $\Delta t_{ijt}$ and $\Delta b_{ijt}$, respectively. These values are time-dependent, as they may vary throughout the day.
[…]
Figure 2: This diagram illustrates a multi-layer network representing an Electric Autonomous Mobility-on-Demand (EAMoD) system. Each layer corresponds to a different battery state of charge (SoC). When a vehicle travels in the network (illustrated by a dashed arrow with a vehicle icon), it consumes energy, leading to a decrease in its state of charge (SoC). In contrast, a dashed arrow with a charging station represents a charging trip, during which the vehicle remains at the station and its SoC increases.
2.2.1 Flow dynamics
There are three types of vehicle movements in the network model. First, passenger transport is represented by $x^{\text{p}}_{ijbt}$, denoting vehicles traveling from origin $i$ to destination $j$ with SoC $b$ at time $t$. Second, vehicle relocation is modeled using $x^{\text{r}}_{ijbt}$, which represents vehicles moving between stations without passengers, a crucial mechanism for balancing vehicle distribution across the network. Third, vehicles engaged in charging or V2G operations are represented by $x^{\text{c}}_{iibt}$ and $x^{\text{d}}_{iibt}$, respectively. The index $ii$ is used for these operations because charging and V2G are modeled as trips within the same station. To ensure conservation of vehicles in the system, we enforce a flow conservation constraint at each station:
$x_{iibt}^{\text{c}}-x_{ii(b-\Delta b^{\text{c}})(t-1)}^{\text{c}}+x_{iibt}^{\text{d}}-x_{ii(b+\Delta d)(t-1)}^{\text{d}}+\sum_{j=1}^{N}\left[x_{ijbt}^{\text{p}}+x_{ijbt}^{\text{r}}-x_{ji(b+\Delta b_{jit})(t-\Delta t_{jit})}^{\text{p}}-x_{ji(b+\Delta b_{jit})(t-\Delta t_{jit})}^{\text{r}}\right]=\phi_{ibt}$ (1)
$\qquad\forall i\in\mathcal{N},\ t\in\mathcal{T},\ \forall b\in\mathcal{B}.$
where $\Delta b_{jit}$ denotes the expected SoC reduction during a trip from station $j$ to station $i$, and $\Delta b^{\text{c}}$ and $\Delta d$ represent energy gained or lost through charging and V2G operations, respectively.
The first four terms in Eq. (LABEL:eq:flowcons) represent vehicles entering or exiting charging or V2G services. The summation terms account for vehicles departing (first two terms) and arriving (last two terms) at each station. This equation ensures system-wide vehicle conservation while incorporating temporal and energy constraints. Specifically, each vehicle movement is subject to a travel time delay $\Delta t_{jit}$ and an associated SoC reduction $\Delta b_{ijt}$, given by:
$\Delta b_{ijt}=\lfloor\frac{e_{ijt}}{e^{\text{b}}}\rceil,$
where $e^{\text{b}}$ is the battery capacity and $e_{ijt}$ is the energy consumption for travel between stations $i$ and $j$ at time $t$.
2.2.2 Imbalance dynamics
EAMoD systems are prone to imbalances caused by spatial and temporal variations in travel demand. To mitigate these imbalances through effective vehicle allocation, it is essential to model the imbalance at each station. We define the imbalance at a station as the difference between the number of passengers awaiting service and the number of available vehicles. Ideally, system operations should maintain this imbalance at zero. However, due to demand fluctuations and operational constraints, achieving a perfect balance is often infeasible. Therefore, we introduce the following constraint, which allows for a positive imbalance:
$\lambda_{ijt}-\sum_{b=b_{ij}^{\min}}^{1}x^{p}_{ijbt}+s_{ij(t-1)}=s_{ijt},\quad\forall i,j\in\mathcal{N},\forall t\in\mathcal{T},$ (2)
where $\lambda_{ijt}$ denotes the travel demand from station $i$ to station $j$ at time $t$, and $x^{p}_{ijbt}$ represents the number of vehicles with a SoC between $b_{ij}^{\min}$ and 1. The parameter $b_{ij}^{\min}$ indicates the minimum SoC required for a vehicle to complete a trip from $i$ to $j$, accounting for the energy consumption $\Delta b_{ijt}$ and a predefined safety margin. The variable $s_{ijt}$ is a slack variable that captures unmet demand, thereby allowing some flexibility in maintaining balance.
2.2.3 Charging Dynamics
Charging dynamics within the network describe how the SoC of vehicles at each charging node evolves. The increase in SoC, denoted as $\Delta b$, during each time interval is calculated by the equation:
$\Delta b=\left\lceil{\frac{P^{\text{c}}(b)\Delta t}{e^{b}}}\right\rceil.$
In this equation, $P^{\text{c}}(b)$ is the SoC-dependent charging power and $\Delta t$ is the discrete time-step length. The operator $\left\lceil\cdot\right\rceil$ rounds $\Delta b$ up to the nearest integer, ensuring discrete SoC bins. Charging power typically exhibits non-linear behavior with SoC rates slow as the battery approaches full charge. We adopt the charging-power curve $P^{\text{c}}(b)$ from [50], which remains approximately constant up to an SoC of $45\%$ and then decreases linearly.
Within each discrete time interval, vehicles at a charging station may begin a charging session, continue a session from a previous interval, or complete their charging. We assume a one-interval, $\Delta t$, charging duration per decision step. This simplification fits with the receding horizon control strategy used in the model, where decisions are regularly updated. By limiting charging sessions to a single interval, the model remains effective and responsive to changes in charging demand and infrastructure availability, without the need to manage the complexities of prolonged charging durations.
The energy depletion of the battery when the vehicle is V2G is given by:
$\Delta d=\left\lceil{\frac{P^{\text{d}}\Delta t}{e^{b}}}\right\rceil.$
where $P^{\text{d}}$ is the power delivered to the grid, which is considered to be constant. Additionally, the availability of chargers at each station is limited, which necessitates a constraint on the total number of vehicles that can charge or await charging at any given time:
$\sum_{b}x_{ibt}^{\text{c}}+x_{ibt}^{\text{d}}\leq\text{k}^{\text{c}}_{it},\quad\forall i\in\mathcal{N},\forall t\in\mathcal{T},$ (3)
where $\text{k}^{\text{c}}_{i}(t)$ represents the total number of charging stations available at each site. This constraint ensures that the model accurately reflects the actual limitations of the charging infrastructure, preventing over-allocation of charging resources.
2.3 Objective Function
We formulate an economic objective function in which each term corresponds to a specific monetary cost or revenue. The goal is to minimize the expected total cost under uncertainty in travel demand, charging-station occupancy, energy consumption, and travel time. The true, unknown, and time-varying probability distributions of these random variables are denoted by $\mathcal{D}_{\lambda}^{*},\ \mathcal{D}_{\text{k}}^{*},\ \mathcal{D}_{\text{b}}^{*},\ \text{and }\mathcal{D}_{\tau}^{*},$ respectively. The uncertainties in energy consumption $\mathcal{D}_{\text{b}}^{*}$ and travel time $\mathcal{D}_{\tau}^{*}$ are handled using chance constraints, while the distributions $\mathcal{D}_{\lambda}^{*}$ and $\mathcal{D}_{\text{k}}^{*}$ enter directly into the expected cost. The multi-stage SMPC objective is:
$J\!\left(x_{ijbt}^{\mathrm{r}},x_{ijbt}^{\mathrm{p}},x_{iibt}^{\mathrm{c}},x_{iibt}^{\mathrm{d}}\right)$ $=\underbrace{\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{b=0}^{1}Q\!\left(x_{ijb0}^{\mathrm{r}},x_{ijb0}^{\mathrm{p}},x_{iib0}^{\mathrm{c}},x_{iib0}^{\mathrm{d}}\right)}_{\text{stage }t=0}$ (4)
$\quad+\mathbb{E}_{\lambda_{1}\sim\mathcal{D}_{\lambda_{1}},\,k^{\text{c}}_{1}\sim\mathcal{D}_{\text{k},1}}\Bigg{[}\mathbb{E}_{\lambda_{2}\sim\mathcal{D}_{\lambda_{2}|\lambda_{1}},\,k^{\text{c}}_{2}\sim\mathcal{D}_{\text{k},2|k^{\text{c}}_{1}}}\Bigg{[}\cdots$
$\qquad\cdots\mathbb{E}_{\lambda_{T-1}\sim\mathcal{D}_{\lambda_{T-1}|\lambda_{1:T-2}},\,k^{\text{c}}_{T-1}\sim\mathcal{D}_{\text{k},T-1|k^{\text{c}}_{1:T-2}}}\Bigg{[}\sum_{t=1}^{T}Q(x_{ijbt}^{\text{r}},x_{ijbt}^{\text{p}},x_{iibt}^{\text{c}},x_{iibt}^{\text{d}})\Bigg{]}\cdots\Bigg{]}\Bigg{]}$
$+\mathbb{E}_{\lambda_{T},k_{T}^{c}\sim(\mathcal{D}_{\lambda}^{*},\mathcal{D}_{k}^{*})}\Bigg{[}\min\!\big{(}q_{\max},\max\!\big{(}(b_{\max}-b)a,0\big{)}\big{)}$
$\qquad\qquad\left(\sum_{i,j,b}x_{ijbT}^{\mathrm{r}}+x_{ijbT}^{\mathrm{p}}+\sum_{i,b}x_{iibT}^{\mathrm{c}}+x_{iibT}^{\mathrm{d}}\right)\Bigg{]},$
which comprises three main components. The first is the stage cost, defined as
$Q(x_{ijbt}^{\text{r}},x_{ijbt}^{\text{p}},x_{iibt}^{\text{c}},x_{iibt}^{\text{d}})=c_{ijb}^{\text{r}}x_{ijbt}^{\text{r}}-c_{ijb}^{\text{p}}x_{ijbt}^{\text{p}}+c_{ibt}^{\text{c}}x_{iibt}^{\text{c}}-c_{ibt}^{\text{d}}x_{iibt}^{\text{d}},$
where the coefficients $c_{ijb}^{\text{r}},c_{ijb}^{\text{p}},c_{ibt}^{\text{c}},$ and $c_{ibt}^{\text{d}}$ are described in Table 2. The initial summation in (4) corresponds to this stage cost at time $t=0$. The second term represents the expected cost over the interval $t\in[1,T]$, where $Q$ captures the costs at each future stage and the random variables $\lambda_{it}$ and $k^{\text{c}}_{it}$ are sampled from their respective distributions. Finally, the third term is a terminal cost that penalizes vehicles whose state of charge (SoC) is below the maximum level $b_{\text{max}}$. This penalty accounts for continued operation of the EAMoD system beyond the optimization horizon. The constant $q_{\text{max}}$ provides an upper bound, while the parameter $a$ converts the battery shortfall $(b_{\text{max}}-b)$ into a monetary penalty per unit of missing energy. This design encourages vehicles to end the horizon sufficiently charged, avoiding unnecessary charging delays while still penalizing low SoC levels in proportion to anticipated future service requirements and electricity prices.
Table 2: Cost and profit components used in the optimization model.
Symbol Description
$c_{ijb}^{\text{r}}$ The cost of rebalancing a vehicle from station $i$ to station $j$ with SoC $b$ at time $t$, representing the expected energy consumption cost.
$c_{ibt}^{\text{c}}$ The cost of fast charging a vehicle at station $i$ at time $t$, where the electricity price may vary depending on the time of day.
$c_{ijb}^{\text{p}}$ The profit from transporting customers from station $i$ to station $j$ with SoC $b$ at time $t$. This profit is calculated based on the distance, with a constant cost per distance throughout the day.
$c_{ibt}^{\text{d}}$ The profit from V2G services provided by a vehicle at station $i$ at time $t$. The profitability is influenced by the varying electricity prices throughout the day.
2.4 The Multi-Stage SMPC Problem
The EAMoD control problem for rebalancing and charging under uncertainty can now be formulated using the model and objective function introduced above. The overall goal is to improve both service quality and operational efficiency. This is achieved by strategically rebalancing vehicles to match spatial demand patterns and by scheduling charging and V2G operations to maintain adequate battery SoC.
Uncertainty in the future evolution of the random variables is addressed using a combined scenario-based and robust approach. Travel demand and charger availability are modeled by a scenario tree that approximates the expected value, with each realization indexed by the superscript $s$. In contrast, uncertainties in energy consumption and travel time are handled through a robust, chance-constrained formulation, since violations in these quantities (e.g., depleted batteries or late arrivals) would directly jeopardize system feasibility and reliability. The resulting optimization problem is updated sequentially as new information about the random variables becomes available. The multi-stage SMPC problem is given by
$\min_{x_{ijbt}^{\text{r},s},\,x_{ijbt}^{\text{p},s},\,x_{iibt}^{\text{c},s},\,x_{iibt}^{\text{d},s}}J_{N}(x_{ijbt}^{\text{r},s},x_{ijbt}^{\text{p},s},x_{iibt}^{\text{c},s},x_{iibt}^{\text{d},s})$ (5a)
subject to
[…]
$\mathbb{P}(\Delta b_{ijt}\leq\beta)\geq 1-\epsilon^{\text{b}},$ (5c)
$\mathbb{P}(\Delta t_{ijt}\leq\tau)\geq 1-\epsilon^{\text{t}},$ (5d)
$x_{ijbt}^{\text{r},s}=x_{ijbt}^{\text{r},s^{\prime}},\quad x_{ijbt}^{\text{p},s}=x_{ijbt}^{\text{p},s^{\prime}},\quad x_{iibt}^{\text{c},s}=x_{iibt}^{\text{c},s^{\prime}},\quad x_{iibt}^{\text{d},s}=x_{iibt}^{\text{d},s^{\prime}}\qquad\forall s,s^{\prime}\in S_{t}:\ (\xi^{s}_{0},...,\xi^{s}_{t-1})=(\xi^{s^{\prime}}_{0},...,\xi^{s^{\prime}}_{t-1}),$ (5e)
$x_{ijbt}^{\text{p}},\ x_{iibt}^{\text{c}},\ x_{iibt}^{\text{d}}\in\mathbb{N}\qquad\forall i,j\in N,\ t\in\mathcal{T},\ b\in\mathcal{B},$ (5f)
$x_{ijbt}^{\text{r}}\in\mathbb{R}^{+}\qquad\forall i,j\in N,\ t\in\mathcal{T},\ b\in\mathcal{B}.$ (5g)
Constraint (5e) enforces non-anticipativity: at each stage, decisions must be identical across all scenarios that share the same history. Constraints (5c) and (5d) are chance constraints requiring that, with probabilities at least $1-\epsilon^{\text{b}}$ and $1-\epsilon^{\text{t}}$, energy consumption and travel time remain below thresholds $\beta$ and $\tau$. Smaller values of $\epsilon$ correspond to more conservative (robust) policies.
The problem is nearly totally unimodular (TU). In general, the LP relaxation of a TU mixed-integer linear program (MILP) yields integer solutions [55]. In our case, however, two constraints break TU. First, Eq. (2) enforces integrality on the sum of passenger-serving variables $x_{ijbt}^{\text{p}}$ rather than on each variable individually. Second, Eq. (3) couples charging and discharging decisions across vehicles in a way that also violates TU. Consequently, explicit integrality must be imposed on $x_{ijbt}^{\text{p}},x_{iibt}^{\text{c}},$ and $x_{iibt}^{\text{d}}$. Only the rebalancing variables $x_{ijbt}^{\text{r}}$ can be treated as positive continuous flows. The overall optimization problem therefore remains a MILP but cannot exploit full TU structure. A significant computational challenge in solving Eq. (5) arises from the exponential growth in the number of scenarios considered at each stage, as also noted by [56]. To address this, the next section introduces an integrated scenario prediction and reduction method, together with a decomposition technique.
3 Scenario Tree Generation and Reduction
To optimize fleet rebalancing and charging under uncertainty, we use a SMPC framework. A key challenge is the uncertain evolution of future travel demand and charger availability, which prevents a closed-form evaluation of the expected cost. We address this by representing uncertainty with a scenario tree that captures both multiple realizations and temporal dependence. This section outlines the construction of the scenario tree, the Bayesian Neural Network (BNN)-based forecasting methodology, the scenario reduction procedure, and the integration of the reduced tree into the SMPC formulation.
3.1 Scenario Tree Construction
The true distributions $\mathcal{D}_{\lambda}^{*}$ and $\mathcal{D}_{\text{k}}^{*}$ for travel demand and charger availability are unknown, so the expectation in (4) cannot be computed directly. To approximate this, we represent uncertainty using a scenario tree, which encodes multiple realizations of the random variables and their evolution over time (see Fig. 3). At each time step $t$, a scenario $s\in\{1,2,\dots,S\}$ is defined as $\xi^{s}_{t}=[\lambda_{ijt}^{s},k_{it}^{c,s}]$, where $\lambda_{ijt}^{s}$ is the predicted travel demand and $k_{it}^{c,s}$ the available chargers in scenario $s$. The cost function is then reformulated using the Sample Average Approximation (SAA) approach:
$\hat{J}_{N}(x_{ijbt}^{\text{r}},x_{ijbt}^{\text{p}},x_{iibt}^{\text{c}},x_{iibt}^{\text{d}})=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{b=0}^{1}Q(x_{ijb0}^{\text{r}},x_{ijb0}^{\text{p}},x_{iib0}^{\text{c}},x_{iib0}^{\text{d}})$ (6)
$+\sum_{t=1}^{T}\sum_{s=1}^{S}p_{s}Q(x_{ijbt}^{\text{r,s}},x_{ijbt}^{\text{p,s}},x_{iibt}^{\text{c,s}},x_{iibt}^{\text{d,s}},\lambda_{ijt}^{s},k_{it}^{\text{c,s}})$
$+p_{s}\min(q_{max},\max((b_{\text{max}}-b)a,0))(x_{ijbT}^{\text{r,s}}+x_{ijbT}^{\text{p,s}}+x_{iibT}^{\text{c,s}}+x_{iibT}^{\text{d,s}}),$
where $p_{s}$ denotes the probability of scenario $s$, and $\sum_{s=1}^{S}p_{s}=1$. The function $\hat{J}_{N}(\cdot)$ approximates the original objective in Eq. (4), where the expected value is replaced with a scenario tree. To avoid exponential growth of the scenario tree with respect to the prediction horizon, we introduce the concept of a robust horizon, beyond which uncertainty is assumed to remain constant. This idea, proposed in [57], allows branching only up to a certain stage, making the tree more tractable (see Fig. 3). Additionally, to further limit unnecessary tree growth when the time discretization is fine, scenarios are assumed to remain constant over short time intervals, effectively reducing the number of unique branches in closely spaced time steps.
[…]
Figure 3: Scenario tree representation of the random variables, incorporating both the robust horizon and prediction horizon. Each scenario is defined as $\xi^{s}_{t}=(\lambda^{s}_{ijt},k_{i,t}^{c,s})$.
3.2 Forecasting Travel Demand and Charger Availability
Given historical spatio-temporal data on travel demand and charger availability, our goal is to predict future values, quantify uncertainty, and model correlations across time and space. Input features $x$ include time, location, and contextual variables such as weather or day of the week. Travel demand is recorded as the number of trips between discretized spatial zones over fixed time intervals, resulting in matrices $\lambda_{ijt}\in\mathbb{Z}^{N\times N}$. Charger availability is given as vectors $k_{it}^{c}\in\mathbb{Z}^{N}$. Because both outputs are non-negative counts, the predictive model must produce integer-valued outputs and capture predictive uncertainty.
We use a Bayesian Neural Network (BNN) to model these forecasts. Unlike standard neural networks, BNNs maintain distributions over their weights $\theta$, allowing for explicit modeling of epistemic uncertainty. Given input features $x$, the neural network defines a parametric function $f(x;\theta)$ that produces an output $\hat{y}$. The Bayesian approach introduces a posterior distribution over the parameters:
$p(\theta|\mathcal{D})\propto p(\mathcal{D}|\theta)p(\theta),$ (7)
where $p(\theta)$ is the prior over weights and $p(\mathcal{D}|\theta)$ is the likelihood of the observed data given parameters $\theta$. Prediction for a new input $x$ involves marginalizing over this posterior:
$p(\hat{y}|x,\mathcal{D})=\int p(\hat{y}|x,\theta)p(\theta|\mathcal{D})d\theta.$ (8)
This integral is intractable in general, so approximate inference techniques such as Variational Inference (VI) or Markov Chain Monte Carlo (MCMC) are used. In our work, we adopt the BNF model proposed in [51], which is specifically designed for spatio-temporal prediction with multiple correlated outputs. It models the predictive distribution using the Negative Binomial (NB) distribution, which is well-suited for overdispersed count data:
$p(\hat{y}(x)|\mathcal{D})\sim\mathrm{NB}(\hat{\mu},\hat{\sigma}^{2}),$
where $\hat{\mu}$ and $\hat{\sigma}^{2}$ are the mean and variance predicted by the model. These predictive distributions form the basis for scenario generation in the next step of our stochastic optimization framework. An important advantage of the Bayesian formulation is that the model can be updated sequentially as new data arrive. In the variational inference setting, this is achieved by using the previously learned posterior as the prior when incorporating a new dataset $\mathcal{D}_{t}$. This enables efficient online learning without retraining from scratch, while preserving past information. Such posterior-to-prior updates allow the BNN to adapt to changing travel demand patterns or charger availability in real time, which is crucial for EAMoD operations. These predictive distributions, updated over time, form the basis for scenario generation in the next step of our stochastic optimization framework.
3.3 Scenario Tree Reduction
From the BNF model, we generate scenario samples of the form:
$\xi^{s}=(\xi_{1}^{s},\xi_{2}^{s},\dots,\xi_{T}^{s}),\quad s=1,\dots,S,$
where each $\xi_{t}^{s}$ consists of sampled travel demand and charger availability values for scenario $s$ at time $t$. This collection of sampled sequences is referred to as a scenario fan. To ensure tractability, we reduce the number of scenarios using clustering techniques. One common approach is to treat the problem as a two-stage stochastic program, but this requires a large number of samples. Instead, we construct a reduced scenario tree that preserves the structure of a multi-stage stochastic process. The tree is built iteratively from the root. At each stage, we sample future values conditionally using the predictive model:
$\xi^{s}_{(t+1)}\sim P(\hat{y}\mid\mathcal{D},\hat{y}(x(t))).$
We then apply soft $k$-means clustering to reduce the number of scenarios, using the cluster centroids as representatives for each scenario. This process is repeated for each time stage. The number of clusters $K_{t}$ at each step is specified by the user in order to balance tractability and representativeness of the uncertainty.
Alternative reduction approaches based on probability metrics, such as the Wasserstein distance or the Nested distance, are not adopted here due to their high computational cost and limited scalability in multi-stage settings. Instead, our method follows the general idea of clustering-based reductions that have been proposed in previous works, where scenarios are grouped and replaced by representative centroids. This allows us to construct reduced scenario trees that remain computationally feasible while still capturing the main patterns of uncertainty. The full reduction algorithm is given in Algorithm 1.
Algorithm 1 Scenario Tree Reduction
1:Input: Initial state $\xi_{0}$, dataset $\mathcal{D}$, prediction model $P(\cdot\mid\mathcal{D})$, time horizon $T$, number of scenarios per stage $K_{t}$
2:Output: Reduced scenario tree
3:Initialize root node with $\xi_{0}$
4:for $t=0$ to T1T-1 do
5:  Sample $M$ scenarios: $\xi^{s}(t+1)\sim P(\hat{y}(x)\mid\mathcal{D},\hat{y}(x(t)))$
6:  Apply soft $k$-means to obtain $K_{t}$ representative scenarios (with $K_{t}$ specified by the user)
7:  Set cluster centroids
8:  Assign transition probabilities from clustering weights
9:end for
10:return Constructed scenario tree
4 Nested Benders Decomposition
Directly solving the multi-stage stochastic problem in Eq. (5) can be computationally expensive, and even intractable, due to the exponential growth of scenarios over the planning horizon. To address this challenge, we employ Nested Benders Decomposition (NBD), an extension of classical Benders decomposition tailored for multi-stage stochastic programs. The key idea is to decompose the original problem into a hierarchy of two-stage problems, one for each stage and scenario, while iteratively exchanging information between them (see Figs. 4 and 5). This approach reduces the computational burden while preserving anticipativity across the scenario tree.
In the NBD framework, the problem is reformulated as a sequence of nested two-stage subproblems, each associated with a specific stage and scenario (Fig. 4). A master problem (MP) coordinates the solution process by integrating Benders cuts derived from the duals of the subproblems (SP) (Fig. 5). These cuts iteratively refine the feasible region of the MP, guiding the solution process toward global optimality while reducing computational effort.
The NBD procedure is executed in a forward–backward iterative scheme, commonly referred to as fast-forward–fast-back sequencing [58]. During the forward pass, the first-stage problem and subsequent stages, up to one stage before the robust horizon, are treated as master problems. In the backward pass, only the first stage remains a master problem, while all subsequent stages are treated as subproblems. To compute valid duals, integer variables in these subproblems must be relaxed to continuous ones.
The iterative process terminates once convergence is achieved, defined as the gap between the subproblem costs and their estimated values in the MP falling below a specified tolerance, or when a maximum number of iterations is reached. Because subproblem integer variables are relaxed, the resulting solutions may not be integer-feasible. This relaxation introduces a duality gap between the relaxed subproblem objective and the true integer-optimal objective, which weakens the tightness of the Benders cuts. Techniques such as Gomory cuts [59] and Lagrangian cuts [60] can yield tighter cuts, but they often significantly slow convergence and are therefore not adopted in our implementation.
[…]
Figure 4: An example of solving a multi-stage SMPC problem by decomposing it into a nested series of two-stage SMPC problems. The figure on the left illustrates the first and second stages of the SMPC problem, while the figure on the right shows a portion of the second and third stages. Both sections can be resolved using Benders Decomposition.
Master ProblemSubproblem 1Subproblem 2Subproblem NTrial solutionBD cutsTrial solutionBD cutsTrial solutionBD cuts
Figure 5: Illustration of nested Benders decomposition. The master problem proposes trial solutions which are passed sequentially to subproblems at different stages. Each subproblem generates Benders cuts based on its solution and returns them to the previous stage, enabling iterative refinement across the scenario tree
Within the NBD framework, multiple MPs are defined, and as the decision process progresses down the scenario tree, SPs are recursively promoted to MPs. To formalize this structure, we present general formulations for both MPs and SPs. Based on the SMPC formulation in Eq. 5 and the reformulated objective function in Eq. 6, the first-stage master problem can be expressed as follows:
$\operatorname*{minimize}_{x_{ijbt}^{\text{r},s},x_{ijbt}^{\text{p},s},x_{iibt}^{\text{c},s},x_{iibt}^{\text{d},s}}J_{N}(x_{ijbt}^{\text{r},s},x_{ijbt}^{\text{p},s},x_{iibt}^{\text{c},s},x_{iibt}^{\text{d},s})+\sum_{s=1}^{S}p_{s}\alpha_{s}$ (9a)
subject to
[…]
$x_{ijbt}^{\text{r},s}=x_{ijbt}^{\text{r},s^{\prime}},\quad x_{ijbt}^{\text{p},s}=x_{ijbt}^{\text{p},s^{\prime}},\quad x_{iibt}^{\text{c},s}=x_{iibt}^{\text{c},s^{\prime}},\quad x_{iibt}^{\text{d},s}=x_{iibt}^{\text{d},s^{\prime}}\qquad\forall s,s^{\prime}\in S_{t}:(\xi^{s}_{0},...,\xi^{s}_{t-1})=(\xi^{s^{\prime}}_{0},...,\xi^{s^{\prime}}_{t-1})$ (9c)
$\mathbb{P}(\Delta b_{ijt}\leq\beta)\geq 1-\epsilon^{\text{b}}$ (9d)
$\mathbb{P}(\Delta t_{ijt}\leq\tau)\geq 1-\epsilon^{\text{t}}$ (9e)
$\alpha_{s}\geq Q^{s}+\sum_{ijb}\pi_{ijbt}^{\text{r},s}(x_{ijbt}^{\text{r},s}-\bar{x}_{ijbt}^{\text{r},s}))+\pi_{ijbt}^{\text{p},s}(x_{ijbt}^{\text{p},s}-\bar{x}_{ijbt}^{\text{p},s}))+\pi_{iibt}^{\text{c},s}(x_{iibt}^{\text{c},s}-\bar{x}_{iibt}^{\text{c},s}))+\pi_{iibt}^{\text{d},s}(x_{iibt}^{\text{d},s}-\bar{x}_{iibt}^{\text{d},s}))$ (9f)
$\qquad\forall s\in S_{t+1}$ (9g)
$x_{ijbt}^{\text{p}}\in\mathbb{N}\qquad\forall i,j\in N,t\in\mathcal{T},\forall b\in\mathcal{B}.$ (9h)
$x_{ijbt}^{\text{r}},x_{iibt}^{\text{c}},x_{iib}^{\text{d}}(t)\in\mathbb{R}^{+}\qquad\forall i,j\in N,t\in\mathcal{T},\forall b\in\mathcal{B},$ (9i)
where $p_{s}$ is the probability of each scenario, $S_{t}$ is the number of scenarios in the stage, and $\alpha_{s}$ is the approximated lower bound of the cost of the subproblem. Constraint (9f) represents the Benders cuts from the subproblems at the next time step, and $S_{t+1}$ denotes the set of all subproblems at time $t+1$. This version of the Benders cuts is referred to as the compact form. In this constraint, $Q^{s}$ represents the cost of each next-stage subproblem, and $\pi$ denotes the dual variables associated with the corresponding decision variables in the MP. The BD cuts generated in these MPs will always be optimality cuts due to the flow conservation constraints in Eq. (LABEL:eq:flowcons). As a result, feasibility cuts are not required. In this implementation, we use multi-cuts, where one Benders cut is added to the master problem for each subproblem. Alternatively, a single-cut approach can be used, where the cuts from each subproblem are aggregated into a single cut. The advantages and disadvantages of both approaches are discussed in Chapter 5 of [61]. It should be noted that whenever an optimization problem is a master problem, integrality is enforced on $x_{ijbt}^{\text{p},s}$. However, when moving up the scenario tree and the MP becomes a subproblem, the problem is relaxed to a linear program to enable the computation of dual variables. The subproblem is then,
$\operatorname*{minimize}_{x_{ijbt}^{\text{r},s},x_{ijbt}^{\text{p},s},x_{iibt}^{\text{c},s},x_{iibt}^{\text{d},s}}J_{N}(x_{ijbt}^{\text{r},s},x_{ijbt}^{\text{p},s},x_{iibt}^{\text{c},s},x_{iibt}^{\text{d},s})+\sum_{s=1}^{S}p_{s}\alpha_{s}$ (10a)
subject to
[…]
$x_{ijbt}^{\text{r},s}=x_{ijbt}^{\text{r},s^{\prime}},\quad x_{ijbt}^{\text{p},s}=x_{ijbt}^{\text{p},s^{\prime}},\quad x_{iibt}^{\text{c},s}=x_{iibt}^{\text{c},s^{\prime}},\quad x_{iibt}^{\text{d},s}=x_{iibt}^{\text{d},s^{\prime}}\qquad\forall s,s^{\prime}\in\mathcal{S}_{t}:(\xi^{s}_{0},...,\xi^{s}_{t-1})=(\xi^{s^{\prime}}_{0},...,\xi^{s^{\prime}}_{t-1})$ (10c)
$\mathbb{P}(\Delta b_{ijt}\leq\beta)\geq 1-\epsilon^{\text{b}}$ (10d)
$\mathbb{P}(\Delta t_{ijt}\leq\tau)\geq 1-\epsilon^{\text{t}}$ (10e)
$\alpha_{s}\geq Q^{s}+\sum_{ijb}\pi_{ijb}^{\text{r},s}(x_{ijb}^{\text{r},s}-\bar{x}_{ijb}^{\text{r},s}))+\pi_{ijb}^{\text{p},s}(x_{ijb}^{\text{p},s}-\bar{x}_{ijb}^{\text{p},s}))+\pi_{iib}^{\text{c},s}(x_{iib}^{\text{c},s}-\bar{x}_{iib}^{\text{c},s}))+\pi_{iib}^{\text{d},s}(x_{iib}^{\text{d},s}-\bar{x}_{iib}^{\text{d},s}))$ (10f)
$\qquad\forall s\in S_{t+1}$ (10g)
$x_{ji(b+\Delta b_{ji})(t-\Delta t_{ji})}^{\text{p}}=\bar{x}_{ji(b+\Delta b_{ji})(t-\Delta t_{ji})}^{\text{p}}$ (10h)
$x_{ji(b+\Delta b_{ji})(t-\Delta t_{ji})}^{\text{r}}=\bar{x}_{ji(b+\Delta b_{ji})(t-\Delta t_{ji})}^{\text{r}}$ (10i)
$x_{ii(b-\Delta b^{c})(t-1)}^{\text{c}}=\bar{x}_{ii(b-\Delta b^{c})(t-1)}^{\text{c}}$ (10j)
$x_{ii(b+\Delta d)(t-1)}^{\text{d}}=\bar{x}_{ii(b+\Delta d)(t-1)}^{\text{d}}$ (10k)
$s_{ij(t-1)}=\bar{s}_{ij(t-1)}$ (10l)
$x_{ijbt}^{\text{p}},x_{ijbt}^{\text{r}},x_{iibt}^{\text{c}},x_{iib}^{\text{d}}(t)\in\mathbb{R}^{+}\qquad\forall i,j\in N,t\in\mathcal{T},\forall b\in\mathcal{B}.$ (10m)
where constraints eqs.˜10j, 10i, 10h, 10k and 10l are the solution from the previous time step MP. The SP have the BD cuts from previous iteration of the NB, eq.˜10f.
4.1 Improving Nested Benders Decomposition
To accelerate the convergence of the NBD algorithm, we employ several enhancement techniques. First, subproblems corresponding to the same stage can be solved in parallel, which can significantly reduce computational time and particularly when stage aggregation is used [62]. Aggregating multiple stages reduces the depth of the decomposition tree and shifts computational effort from cut generation to problem solving. While solving each aggregated subproblem becomes more time-consuming, the overall number of Benders cuts required typically decreases, leading to faster convergence.
So-called warm-up cuts are implemented prior to starting the NBD algorithm. These cuts are based on simple heuristics. For each scenario, the vehicle imbalance at every station and time step is computed in the absence of rebalancing. Based on these imbalances, slack-penalizing constraints are added to the master problems to incentivize early rebalancing and charging actions. The cost of the slack variable corresponds to the lost revenue from not serving a customer. These warm-up cuts help guide the initial iterations toward more balanced solutions and can significantly reduce convergence time.
5 Mobility network simulations and sensitivity analysis
We evaluate our approach in two high-fidelity case studies, San Francisco and Chicago, constructed from large-scale taxi trip records. We first assess the accuracy of the proposed travel-demand predictions on these data. Next, we analyze the computational performance of the Nested Benders Decomposition (NBD) used to solve the underlying stochastic optimization problems. We then benchmark the proposed stochastic and robust algorithm in the AMoDeus simulator [52]. Finally, we perform a sensitivity analysis to quantify how vehicle energy efficiency, battery capacity, fleet size, and charger availability affect system performance.
5.1 Simulation environment
We evaluate our proposed algorithm using the high-fidelity transport simulator AMoDeus [52], an extension of the Multi-Agent Transport Simulation framework (MATSim) [63] specifically designed for AMoD systems. Two urban contexts are studied: San Francisco and Chicago, both based on real-world taxi data.
The San Francisco case study utilizes taxi data from 2008, encompassing 464,045 customer trips collected from 500 taxis [64]. For simulation, we selected Thursday, May 29, 2008 with a total travel demand of 11,453 trips and an average trip distance of 3.31 km. The Chicago scenario is based on taxi data collected between 2013 and 2023, representing over 212 million trips [65]. Our analysis focuses on Wednesday, July 24, 2019 which recorded a total travel demand of 32,406 trips and an average trip distance of 14.64 km, using data from July 15–24, 2019. To meet the respective demand levels, we simulate fleets of 450 vehicles in San Francisco and 1,400 in Chicago.
In both cities, vehicles are initially distributed across $\mathcal{N}=10$ charging stations, each starting at 80% SoC. The SoC is discretized into $\mathcal{B}=50$ intervals. The simulation horizon is $T=20$ time steps (200 minutes) with a time step length of $\Delta t=10$ minutes, and operational planning is carried out over a robust horizon of 6 time steps (60 minutes). Vehicles are required to end the simulation with at least $b_{\text{end}}=30\%$ SoC. An overview of all simulation parameters is provided in Table 3.
Vehicle battery capacities are set to $e^{b}=40$ kWh for San Francisco and $e^{b}=70$ kWh for Chicago. We use a non-linear charging curve derived from [50], in which charging proceeds at a constant 100 kW until 45% SoC and then decreases linearly to 30 kW at 95% SoC. The stochastic optimization model applies scenario reduction per time step with levels [3, 3, 5, 5, 5], resulting in 1,125 total scenarios. Battery violation tolerance is set to $\epsilon^{b}=0.2$ and demand violation tolerance to $\epsilon^{t}=0.2$. Nested Benders Decomposition (NBD) is used with a gap tolerance of $10^{-2}$ and a maximum of 30 iterations.
A key enhancement to the simulation environment is the integration of an energy consumption estimation module, as AMoDeus does not include this functionality for transportation simulations. This module estimates energy consumption using detailed speed profiles based on the established model from [66]. Furthermore, to accurately represent real-world driving conditions, stochasticity was introduced into the simulation by modeling traffic light behavior and randomizing vehicle speeds.
For electricity pricing, we developed three distinct scenarios (Fig.6): a flat-price profile, a peak-price profile, and a solar-price profile reflecting midday price drops from high solar generation, with the first two derived from Nordpool data [67]. These represent daily operational variations rather than annual averages. For benchmarking the proposed algorithm, vehicle type vehicle 1 is used, while the sensitivity analysis compares three different vehicle types with varying energy consumption rates (Fig.8), under the peak-price scenario only.
All simulations were executed on a MacBook Pro with a 2.3 GHz Quad-Core Intel Core i7 processor and 16 GB of RAM. Optimization problems were solved using the Gurobi optimization library [68].
Figure 6: Electricity prices in €/kWh for three different pricing scenarios: Flat, Peak, and Solar Price, shown over a 24-hour period.
Table 3: Simulation Parameters Overview
Common Parameters
Parameter Value
Number of stations ($\mathcal{N}$) 10
Simulation horizon ($T$) 20 time steps (200 minutes)
Robust horizon 6 time steps (60 minutes)
Scenarios reduction per time step [3,3,5,5,5] (1125 scenarios in total)
Time step length ($\Delta t$) 10 minutes
Number of SoC intervals ($\mathcal{B}$) 50
Minimum required SoC at end ($b_{\text{end}}$) 30 %
V2G power ($P^{d}$) 22 kW
Battery violation tolerance ($\epsilon^{b}$) 0.2
Demand violation tolerance ($\epsilon^{t}$) 0.2
Vehicle type V1
NBD gap tolerance $10^{-2}$
Max number of BD iteration 30
City-Specific Parameters
Parameter San Francisco Chicago
Number of vehicles 450 1400
Battery capacity ($e^{b}$) 40 kWh 70 kWh
Total number of chargers 100 200
Total demand 11,453 32,406
Simulation date May 29, 2008 July 24, 2019
5.2 Travel Demand Prediction
We conducted a comprehensive comparative analysis between BNF, our proposed prediction framework, and three established baselines: multi-output Gaussian Process regression with a spectral kernel (MOSK) [69], Long Short-Term Memory (LSTM) networks, and Vector AutoRegressive Moving Average (VARIMA) models [70]. These baselines were selected based on their proven effectiveness in travel demand forecasting and represent a diverse set of statistical and machine learning approaches [71, 18]. Like BNF, MOSK provides explicit uncertainty estimates, enabling a direct comparison of probabilistic forecasting performance. All models were evaluated on real-world datasets from San Francisco and Chicago. The BNF architecture used two hidden layers with 64 units each. Input features included time, origin coordinates (latitude and longitude), and destination coordinates. A Negative Binomial distribution was used as the prior predictive model to reflect the count nature of the data.
Model performance was assessed using two key metrics which were Root Mean Square Error (RMSE) and Mean Interval Score (MIS). RMSE quantifies the average deviation between predicted and actual values, with lower values indicating higher accuracy. MIS penalizes both the width and coverage of prediction intervals, providing a measure of uncertainty calibration; lower values correspond to sharper and more reliable probabilistic forecasts. Results are shown in Table 4.
BNF achieved the best performance across all scenarios in both RMSE and MIS. In San Francisco, BNF reduced RMSE significantly by an average of 36.6% compared to LSTM, 29.7% compared to VARIMA, and 19.7% compared to MOSK. The improvements were more pronounced in Chicago, with average RMSE reductions of 66.7%, 47.1%, and 15.8% compared to LSTM, VARIMA, and MOSK, respectively. For uncertainty quantification, BNF also demonstrated a clear advantage, reducing MIS by an average of 42.7% in San Francisco and 35.5% in Chicago relative to MOSK.
In addition to demonstrating strong average performance, BNF showed greater robustness during peak demand periods (06:00–10:00 and 17:00–21:00), which are especially important for real-world EAMoD operations. During these high-demand windows, models such as MOSK exhibited significantly higher errors. For example, during the evening peak in Chicago, BNF reduced RMSE by 49.2% compared to MOSK, indicating its superior ability to handle sharp fluctuations in travel demand. The substantial reductions in both RMSE and MIS contribute to more reliable scenario trees. A lower RMSE results in a more accurate average scenario, while a lower MIS improves the estimation of scenario probabilities within the scenario tree.
As shown in Fig. 7, BNF produces accurate travel demand forecasts with well-calibrated 95% prediction intervals at selected Chicago pickup locations. The model effectively captures both the temporal dynamics of demand and the underlying predictive uncertainty.
Table 4: Comparison of RMSE and MIS for different forecasting models in San Francisco and Chicago across key time periods.
Metric Algorithm San Francisco Chicago
06:00–10:00 10:00–14:00 17:00–21:00 06:00–10:00 10:00–14:00 17:00–21:00
RMSE LSTM 14.5175 18.8251 24.2956 27.4708 85.3836 133.4684
VARIMA 18.0840 9.6063 12.0309 29.3161 26.6069 35.5457
MOSK 10.0874 7.8890 10.5079 26.0935 26.0016 28.8127
BNF 8.1455 6.9751 8.0259 25.2362 13.6507 14.6357
MIS MOSK 83.2670 63.8836 77.2634 193.0978 154.8194 148.6521
BNF 40.0050 33.6349 53.0234 110.9026 98.2883 103.8081
Figure 7: Predicted travel demand with 95% prediction intervals for selected locations. Black dots show training data, white dots indicate test data, the red line is the median prediction, and blue shaded areas represent the 95% prediction intervals.
5.3 Nested Benders Decomposition Performance
We analyze the computational performance of Nested Benders Decomposition (NBD), which is used to solve the multistage stochastic optimization problems underlying our SMPC approach. We do not compare against the extensive form of the stochastic problem, as it becomes computationally intractable beyond very small instances. Instead, we evaluate NBD in terms of runtime, convergence behavior, and scalability with respect to the number of scenarios.
Table 5 reports the runtime, number of iterations, and the relative optimality gaps across three scenario set sizes. For the smallest instance (135 scenarios), NBD converges within 10 iterations and achieves a near-exact solution with a final gap of 0.08%. Increasing the number of scenarios to 540 raises the runtime to about 90 seconds, but convergence is still reached in fewer than 10 iterations, with a very small median gap of 0.09%.
For the largest instance (1035 scenarios), the runtime increases to nearly 6 minutes and the number of iterations to 15. While the maximum gap observed during the iterations is higher (4.52%), the algorithm ultimately converges to a high-quality solution with a median gap of 1.50%. These results show that runtime grows with the number of scenarios, but NBD remains tractable and consistently produces solutions with small gaps even for problem sizes where the extensive form would be unsolvable.
Table 5: Runtime and convergence diagnostics across scenario counts.
Scenarios Runtime (sec) Iterations Max Gap (%) Median Gap (%)
135 17.15 10 0.08 0.08
540 90.26 9 0.79 0.09
1035 355.55 15 4.52 1.50
5.4 EAMoD Optimization Performance Comparison
We benchmark our proposed SMPC algorithm, which explicitly accounts for uncertainty in predictions, against three alternative control strategies:
  • Deterministic Mean Prediction Optimization (DMP): Uses only the expected values of predicted travel demand, charging station occupancy, travel time, and energy consumption from BNF, without considering uncertainty.
  • Robust Chance-Constrained Optimization (RCC): A robust approach inspired by [12], designed to guarantee performance under worst-case prediction errors.
  • Reactive Baseline (RB): A reactive strategy with no forecasting. Vehicles are charged whenever the state of charge (SoC) falls below 20%, up to 80%. No vehicle-to-grid (V2G) operations are considered.
All controllers are re-optimized every 10 minutes using a receding horizon approach. The results for Chicago and San Francisco are summarized in Tables  6 and  7.
In Chicago, SMPC achieves the lowest passenger delays. The median waiting time is 214 s, essentially the same as for DMP (215 s), but significantly lower at the 95th percentile with 767 s compared to 872 s for DMP (12% improvement), 822 s for RCC, and 955 s for the reactive baseline (20% improvement). Operationally, SMPC requires 70.1 km of rebalancing per vehicle, which is 27% less than RCC (96.6 km), while keeping the total distance traveled (500.7 km) close to RCC (510.1 km). RB travels the least overall distance (456.1 km), but this comes at the cost of much longer wait times.
In San Francisco, SMPC again dominates in terms of service quality. The median waiting time is 229 s, compared to 240 s for DMP (5% higher), 254 s for RCC, and 360 s for the reactive baseline, corresponding to a 36% improvement over RB. At the 95th percentile, SMPC records 548 s compared to 584 s for DMP, 559 s for RCC, and 670 s for RB, improving extreme waits by almost 20% relative to the reactive baseline. Operationally, SMPC requires 108.5 km of rebalancing per vehicle compared to 124.8 km for RCC, a 13% reduction, while keeping total distance traveled (414.1 km) lower than RCC (431.6 km). Although DMP uses slightly less rebalancing (99.5 km), SMPC achieves consistently lower waiting times across the distribution, showing that the modest increase in repositioning yields substantially improved service reliability.
Energy costs and V2G revenues show similar trends in both cities. Under flat tariffs, all methods incur costs of around €0.06/kWh with negligible revenue. However, under peak and solar tariffs, SMPC and DMP consistently achieve the lowest effective energy costs and the highest revenues. In Chicago, SMPC earns almost €975 per vehicle under solar pricing, which is approximately double the value in San Francisco of €509, reflecting the larger fleet and heavier charging activity. RCC captures somewhat less revenue due to its conservative scheduling, while RB never participates in V2G and pays the highest electricity prices.
Overall, these results demonstrate that the SMPC model strikes a balance between service quality, operational efficiency and energy costs. It delivers shorter wait times than deterministic or reactive baselines, avoids the conservatism of robust optimization and captures meaningful grid revenues by exploiting dynamic electricity prices.
Table 6: Operational performance in Chicago (fleet size 1600) and San Francisco (fleet size 500) under four control strategies. Metrics include mean, median (50% quantile), and 95% quantile wait times (WT), as well as total and rebalance distance per vehicle.
City Method Median 95% Total Dist./Veh. Reb. Dist./Veh.
$\mathrm{s}$] $\mathrm{s}$] $\mathrm{k}\mathrm{m}$] $\mathrm{k}\mathrm{m}$]
Chicago RB 272 955 456.1 4.6
DMP 215 872 490.0 65.1
RCC 220 822 510.1 96.6
SMPC 214 767 500.7 70.1
San Francisco RB 360 670 401.1 3.8
DMP 240 584 407.1 99.5
RCC 254 559 431.6 124.8
SMPC 229 548 414.1 108.5
Table 7: Average electricity price paid per kilowatt-hour (€/kWh) and corresponding vehicle-to-grid (V2G) revenue (€) under flat, peak, and solar pricing schemes for Chicago and San Francisco.
City Method Flat Price Flat V2G Rev. Peak Price Peak V2G Rev. Solar Price Solar V2G Rev.
[€/kWh] [€] [€/kWh] [€] [€/kWh] [€]
Chicago RB 0.06 0.0 0.083 0.00 0.09 0.00
DMP 0.06 0.0 0.078 565 0.07 981
RCC 0.06 0.0 0.076 547 0.08 944
SMPC 0.06 0.0 0.075 563 0.07 975
San Francisco RB 0.06 0.0 0.08 0.00 0.090 0.00
DMP 0.06 0.0 0.07 290 0.070 515
RCC 0.06 0.0 0.08 285 0.074 453
SMPC 0.06 0.0 0.069 291 0.068 509
5.5 Experimental setup for sensitivity analysis
This section investigates the impact of battery size and vehicle energy efficiency on the performance of urban EAMoD systems in San Francisco and Chicago. We analyze three distinct vehicle types, representing high, medium, and low energy efficiency, and consider three battery capacities: 40, 70, and 100 kWh (Table 8). Vehicle parameters, including those determining efficiency, are summarized in Table 9, based on the energy consumption model from [66]. Vehicle weights listed in the table exclude battery mass, which is added separately assuming a specific energy of 268 Wh/kg [72]. Fig.8 illustrates the energy consumption rates of the three vehicle types with a 40 kWh battery, excluding auxiliary loads.
Table 8: Overview of the experimental setup for the mobility-on-demand simulations in San Francisco and Chicago. The different vehicle types corresponds to different energy efficiency (low, medium and high).
Factor San Francisco Chicago
Vehicle Type 1 (low), 2 (medium), 3 (high) 1 (low), 2 (medium), 3 (high)
Battery Size (kWh) 40, 70, 100 40, 70, 100
Fleet Size (vehicles) 450, 500, 550 1400, 1600, 1800
Number of Chargers 100, 150, 200 200, 300, 400
Table 9: Vehicle parameters used to model energy consumption. Symbols, descriptions, and units are listed in the first three columns.
Symbol Description Unit Vehicle 1 Vehicle 2 Vehicle 3
$m$ Weight kg 1500 1700 1900
$\eta^{+}$ Efficiency (acceleration) 0.90 0.85 0.80
$\eta^{-}$ Efficiency (deceleration) 0.90 0.85 0.80
$\eta^{\text{nomial}}$ Efficiency (constant speed) 0.90 0.85 0.80
$c_{r}$ Rolling resistance coefficient 0.0072 0.009 0.0108
$c_{d}$ Drag coefficient 0.22 0.25 0.30
$p_{\text{aux}}$ Auxiliary power W 800 900 1000
$p_{\text{auxidle}}$ Auxiliary idle power W 100 100 100
AA Frontal area m2 2.2 2.5 2.8
Figure 8: Energy consumption of the different vehicles without auxiliary power and with a 40 kWh battery.
5.6 Sensitivity analysis of battery size and energy efficiency (San Francisco)
Battery capacity significantly influences system performance, though the extent depends on vehicle efficiency (Table 10). For Vehicle 1, the most energy-efficient option, wait times remain nearly constant across all battery sizes. Increasing the battery from 40 kWh (median 220s) to 100 kWh (median 219s) yields virtually no change in performance. Similarly, for vehicle 1 the 75th and 95th percentile wait times remain below 450s across all capacities. This indicates that Vehicle 1’s service quality is not constrained by battery size, even with the smallest pack.
Vehicles 2 and 3 display greater sensitivity to battery size. For Vehicle 2, the median wait time decreases slightly from 225s at 40kWh to 217s at 70 kWh, while the share of passengers waiting more than 600s falls from 154 to 151. Expanding to 100 kWh provides only marginal change: median 219 s (slightly higher), with >600 s falling to 144. This suggests that a 70 kWh battery is sufficient to mitigate charging as a performance bottleneck for Vehicle 2, with further increases offering diminishing returns. For Vehicle 3, the least efficient option, the effect of battery size is far more pronounced. With 40 kWh, the system exhibits very poor performance, with a median wait time of 325s, a 95th percentile of 2625s, and more than 3600 passengers experiencing delays over 600s. Upgrading to 70 kWh brings dramatic improvements: the median drops to 219s, the 95th percentile to 432s, and delays over 600s fall by more than 95% to just 145. Increasing to 100 kWh does not significantly improve service further, with a median of 222s and 149 delays over 600s. This highlights that for inefficient vehicles, battery size is a critical driver of feasibility, with performance plateauing around 70 kWh.
These results yield two key insights. First, energy-efficient vehicles like Vehicle 1 exhibit lower sensitivity to battery size, enabling effective performance with smaller, lighter battery packs. This directly contributes to overall system efficiency and potentially lower operational costs. Second, while less efficient vehicles require larger batteries to achieve acceptable service quality, the performance gains tend to plateau around 70 kWh. Importantly, these service dynamics also shape revenues from V2G services. Efficient vehicles consistently generate higher V2G revenues, since their lower consumption leaves more capacity for grid interaction. For example, Vehicle 1 earns €304 per vehicle at 40 kWh and rises to €390 at 100 kWh, the highest across all types. Vehicle 3, by contrast, produces the lowest revenues (e.g., €260 at 40 kWh, €360 at 100 kWh), as its higher driving demand limits grid availability. Larger batteries generally boost revenues across all vehicles, but the marginal gains diminish at the upper end: Vehicle 2 increases from €285 at 40 kWh to €342 at 70kWh, while the additional jump to €386 at 100 kWh is relatively small. Thus, the same diminishing returns seen in wait times and charging behavior are mirrored in V2G participation.
These findings are consistently reflected in other system-level metrics. Figures 9 demonstrate that increasing fleet size and charger availability effectively reduce wait times, particularly for the less efficient Vehicles 2 and 3. Vehicle 1, however, continues to exhibit stable performance, characterized by lower 75th percentile and minimal sensitivity to infrastructure constraints. Furthermore, Vehicle 3 has the highest energy consumption and consequently the greatest charging demand, especially during peak hours, see Fig. 10(a). Conversely, Vehicle 1 maintains a smoother, more aligned charging-consumption profile throughout the day.
Fig.11(a) reports the estimated number of daily charge cycles (defined as charging from 20% to 80% state of charge) across vehicle types and battery sizes. As anticipated, larger batteries consistently require fewer charging cycles. Vehicle 1 consistently exhibits the lowest and most stable charging frequency, requiring just over one cycle per day even with a 40 kWh battery. In contrast, Vehicles 2 and 3 demonstrate higher and more variable charging frequencies, particularly with smaller battery configurations. For instance, Vehicle 3 with a 40 kWh pack exhibits a median of 1.5 cycles per day and substantial variability in its charging behavior. Increasing the battery capacity to 70 kWh or 100 kWh nearly halves the median number of cycles and significantly narrows the distribution of charging events.
These observed differences in charging frequency directly correlate with wait time performance. Frequent charging, especially for Vehicle 3 with small batteries, leads to service delays as vehicles become temporarily unavailable for passenger service. Larger batteries effectively mitigate both the number of charge cycles and the occurrence of such extended delays.
In summary, this analysis highlights that efficient vehicles can deliver consistent service with smaller batteries, while less efficient vehicles necessitate larger battery packs to avoid degraded performance. Reducing charge cycle frequency is paramount not only for prolonging battery lifespan but also for ensuring stable and responsive EAMoD service. These results underscore that optimal vehicle selection for urban EAMoD systems must consider not only immediate performance metrics but also long-term operational sustainability. Ultimately, if a vehicle is sufficiently energy efficient, a 40 kWh battery offers a compelling trade-off between cost, performance, and battery longevity.
Table 10: Summary of wait time and per-vehicle distance by battery capacity and vehicle type for San Francisco.
Vehicle SoC Median 75% 95% Wait > 600s TotDist/Veh RebalDist/Veh V2G Total
[kWh] [s] [s] [s] [km] [km] [€]
1 40 220 301 440 137 240 79 304
2 40 225 309 452 154 206 42 285
3 40 325 1129 2625 3643 216 37 260
1 70 219 296 432 145 284 123 347
2 70 217 296 434 151 241 80 342
3 70 219 299 432 145 209 47 316
1 100 219 303 443 134 291 129 390
2 100 219 303 447 144 292 130 386
3 100 222 306 451 149 190 32 360
(a) Median (circle) and 75th percentile (triangle) of wait times across fleet size
(b) Median (circle) and 75th percentile (triangle) of wait times across number of chargers
Figure 9: Comparison of wait time metrics in San Francisco. Fig. a shows metrics across different fleet sizes and vehicle types with 100 chargers. Fig. b shows metrics across different number of charger and vehicle types with a fleet size of 450 vehicles. All plots assume a 40 kWh battery capacity. For clarity, 75th percentile wait times for Vehicle 3 are not shown in the plots, but their values are as follows: for varying fleet size (450, 500, 550 vehicles), the Q75 values are 1129 s, 610 s, and 590 s, respectively; for varying number of chargers (100, 150, 200), the Q75 values are 1129 s, 654 s, and 650 s, respectively.
Figure 10: Comparison of hourly energy charged and consumed across three different vehicle types with 40 kWh battery capacity for (a) San Francisco and 70 kWh for (b) Chicago. Each line represents the total energy (in kWh) aggregated across all vehicles in a fleet. Charging events are shown with dashed lines, and consumption with solid lines.
Figure 11: Observed charge cycles (20–80% Depth of Discharge) for different vehicle types (1, 2, 3) and battery sizes (40, 70, 100 kWh) in (a) San Francisco and (b) Chicago. These distributions highlight the combined influence of vehicle characteristics and geographical location on battery charging patterns.
5.7 Sensitivity analysis of battery size and energy efficiency (Chicago)
We now turn to the results for Chicago to evaluate how battery size and vehicle energy efficiency impact system performance in a different urban context. As in the San Francisco analysis, we examine three vehicle types with varying efficiencies and battery capacities of 40, 70, and 100 kWh. However, Vehicle 3 with a 40 kWh battery is excluded from the analysis, as its limited capacity is insufficient for viable operation in the Chicago setting.
The energy demand patterns in Chicago are more concentrated during peak periods compared to San Francisco, as seen in Fig.10(b). Vehicles with lower efficiency (i.e., Vehicle 3) show a pronounced mismatch between energy consumption and charging, resulting in steep late-day charging requirements. Vehicle 1 again exhibits the most balanced charging-consumption profile, with relatively low and stable energy use throughout the day. This greater stability also enables Vehicle 1 to contribute more consistently to V2G services, particularly with larger battery sizes.
Battery size significantly influences system performance in Chicago, highlighting the high sensitivity of less efficient vehicles to energy storage capacity in high-demand environments. As shown in Table 11, Vehicle 1’s performance remains robust across battery sizes, with consistently low wait times even in the high-demand Chicago setting. The median wait time improves modestly from 214s with a 40kWh battery to 203s at both 70 and 100 kWh. However, a more pronounced improvement is seen in the upper percentiles. The 75th percentile drops from 329s to 258s between 40 and 70kWh and remains unchanged at 100kWh, suggesting that moderate battery increases help reduce occasional peak delays. The 95th percentile shows a particularly notable decline, from 1302s at 40kWh to 673s at both 70 and 100 kWh, cutting the longest delays by nearly half. Similarly, the number of wait instances exceeding 600 s is more than halved, dropping from 4392 to approximately 1900. These upper-percentile improvements highlight that while the median performance of Vehicle 1 is already strong, increasing battery capacity primarily enhances service reliability by mitigating infrequent but severe delays. At the same time, Vehicle 1 delivers the highest V2G revenues, ranging from €572 at 40kWh to €580 at 100 kWh, confirming that efficiency and moderate battery sizing jointly support both passenger service and grid interaction.
In contrast, Vehicle 2 exhibits a much stronger dependency on battery size, with substantial gains in performance as capacity increases. The median wait time drops from 293s at 40 kWh to 221s at 70 kWh, and further to 204s at 100 kWh. Improvements are even more pronounced at the tail: the 75th percentile shrinks dramatically from 2129s at 40 kWh to 426s at 70 kWh and just 266s at 100 kWh. The 95th percentile follows a similar trend, falling from 9772s to 2919s and then to only 692s—indicating a transformation from highly unstable to near-stable performance. Likewise, the number of wait instances exceeding 600 s plummets from 12,176 at 40 kWh to 6606 at 70 kWh and just 2004 at 100 kWh. These improvements confirm that Vehicle 2, while more efficient than Vehicle 3, remains highly sensitive to energy availability and benefits significantly from battery upgrades. However, unlike Vehicle 1, its V2G revenues remain nearly flat across capacities (€563 at 40–70 kWh, €562 at 100 kWh), suggesting that most of its battery energy is absorbed by travel demand rather than enabling grid services.
Vehicle 3, the least efficient of the fleet, shows the greatest sensitivity to battery capacity and consistently struggles under the high-demand Chicago conditions. At 70kWh, its median wait time is 316s, but the upper tail reveals severe congestion: the 75th percentile reaches 5652s, and the 95th percentile exceeds 18,000s. Over 13,000 trips experience delays longer than 10 minutes. Increasing the battery to 100kWh brings some relief—reducing the median to 237s, the 75th percentile to 1072s, and the 95th percentile to 10,830s. The number of long wait times drops to 9912, but remains unacceptably high. These results underscore the fundamental limitations of low-efficiency vehicles operating in dense urban settings, where even substantial battery investments cannot fully overcome their performance bottlenecks. For Vehicle 3, both battery capacity and extensive infrastructure support are essential, yet still insufficient to match the reliability of more efficient alternatives. Consistent with this, Vehicle 3 also generates the lowest V2G revenues (€559 at 70 kWh and €550 at 100 kWh), reflecting that its high consumption leaves little surplus for grid participation and that larger batteries are used primarily to sustain basic service rather than enable additional revenue streams.
Fig. 12 further highlight how infrastructure scale affects performance. As fleet size increases from 1400 to 1800, all vehicles benefit, but the reduction is especially marked for Vehicles 2 and 3. For example, Vehicle 3 sees median wait time drop from over 300 s to around 240 s. Similarly, increasing the number of chargers from 200 to 400 lowers both median wait times and the 75th percentile. These improvements are again more significant for less efficient vehicles, underscoring their greater dependence on robust infrastructure.
Charging frequency data (Fig.11(b)) shows similar trends as in San Francisco. Vehicle 1 requires just over one cycle per day even with a 40 kWh pack, while Vehicles 2 and 3 exhibit more frequent and variable charging, particularly at smaller battery sizes. For instance, Vehicle 2 with a 40 kWh battery has a median of about 1.2 cycles/day, whereas at 70 kWh this drops to around 0.7. Vehicle 3 consistently shows the highest cycle counts, reflecting its greater energy demand and need for frequent recharging. These frequent cycles not only limit availability for passenger trips but also reduce the opportunities for stable V2G participation.
Overall, the Chicago results reinforce the conclusions from San Francisco while highlighting even sharper contrasts due to higher demand intensity. Vehicle 1 remains robust across battery sizes and system configurations, while Vehicles 2 and 3 require both larger batteries and increased infrastructure to maintain acceptable service levels. The results emphasize that pairing efficient vehicles with moderate battery capacities can yield strong, scalable performance, whereas less efficient platforms demand a combination of capacity and infrastructure investment to avoid service degradation. From a V2G perspective, Vehicle 1 again stands out as the most attractive option, combining stable service with the highest and most consistent revenues. Vehicles 2 and 3, by contrast, show limited or declining V2G potential as their capacity is consumed by travel demand, underscoring that efficiency is the primary driver of grid-service opportunities in high-demand urban settings.
Table 11: Summary of wait time and per-vehicle distance by battery capacity and vehicle type for Chicago.
Vehicle SoC Median 75% 95% Wait >600s TotDist/Veh RebalDist/Veh V2G Total
[kWh] [s] [s] [s] [km] [km] [€]
1 40 214 329 1302 4392 512.3 67.1 572
2 40 293 2129 9772 12176 461.7 52.1 563
1 70 203 258 673 1911 500.7 70.1 563
2 70 221 426 2919 6606 510.0 60.6 563
3 70 316 5652 18175 13113 432.6 50.9 559
1 100 203 258 673 1941 492.0 63.5 580
2 100 204 266 692 2004 504.2 68.8 562
3 100 237 1072 10830 9912 466.4 52.8 550
(a) Median (square) and 75th percentile (triangle) of wait times across number of chargers
(b) Median (square) and 75th percentile (triangle) of wait times across number of chargers
Figure 12: Comparison of wait time metrics in Chicago. Fig. a shows metrics across different fleet sizes and vehicle types with 200 chargers. Fig. b shows metrics across different number of charger and vehicle types with a fleet size of 1400 vehicles. All plots assume a 70 kWh battery capacity. For clarity, 75th percentile wait times for Vehicle 3 are not shown in the plots, but their values are as follows: for varying fleet size (1400, 1600, 1800 vehicles), the Q75 values are 5651 s, 2180 s, and 969 s, respectively; for varying number of chargers (200, 300, 400), the Q75 values are 5651 s, 1172 s, and 932 s, respectively.
6 Conclusion
This paper introduced a combined stochastic and robust optimization framework for Electric Autonomous Mobility-on-Demand (EAMoD) operations under multiple uncertainties. By integrating Bayesian neural networks for demand and charger availability forecasting, clustering-based scenario tree reduction, and a tailored Nested Benders Decomposition algorithm, we developed a scalable stochastic model predictive control (SMPC) approach capable of handling temporally correlated, multivariate uncertainties. High-fidelity simulations in San Francisco and Chicago demonstrate that SMPC reduces median passenger wait times by up to 36%, 95th percentile wait times by nearly 20%, rebalancing distance by up to 27%, and electricity costs by more than 35% relative to established baselines. These gains are consistent across two cities with very different demand intensities, spatial layouts, and fleet sizes, underscoring the framework’s robustness and transferability to diverse urban contexts.
Sensitivity analyses revealed that vehicle efficiency and battery size critically affect system performance. Energy-efficient vehicles maintain stable service levels even with small (40 kWh) batteries, requiring fewer charge cycles and exhibiting minimal sensitivity to infrastructure constraints. In contrast, less efficient vehicles show significant performance degradation with small batteries and benefit most from increases up to 70 kWh, beyond which improvements plateau, especially in high-demand settings such as Chicago. Moreover, efficiency not only drives service quality but also expands opportunities for energy-market participation: efficient vehicles consistently generate higher vehicle-to-grid (V2G) revenues, as their lower energy consumption leaves more capacity for grid interaction. Taken together, the results highlight the importance of jointly optimizing control strategies, vehicle selection, and battery configuration to ensure scalable, reliable, and sustainable EAMoD operations.
7 Acknowledgement
This work was co-funded by Vinnova, Sweden through the project: Simulation, analysis and modeling of future efficient traffic systems. This work was in part supported by the Transport Area of Advance within Chalmers University of Technology.
8 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT and DeepL exclusively to improve the readability and language of the manuscript. The scientific content, ideas, and conclusions are entirely the authors’ own. After using these tools, the authors carefully reviewed and edited the text, and they take full responsibility for all content of the published article.
[…]